OpenStudy (anonymous):

If 3x + y = 18, then 2(3x +y) =__________

OpenStudy (anonymous):

treat "3x + y" as one variable, for example v if v = 18, 2v will be twice that, right? therefore, 2(3x+y) = 36 if x is 5, then 2x + 5 = 2(5) + 5 = 15 15 - x = 15 - 5, so 2x + 5 is 10 more than x same principle on 3 as 1, only you're using the distribution property as well... 3x + 3y = 3(x+y) therefore, 3x + 3y = 36 (which is 3*12, since x+y = 12) 3.6 + 6.4 = 10 x + y = 10 100*(x+y) = 100*10, = 1000